Two vectors of magnitude F and 2F have a resultant 1.5F. Calculate the angle between the vectors
1
Expert's answer
2020-07-21T12:39:05-0400
"(1.5F)^2=F^2+(2F)^2+2F(2F)\cos{\theta}"

"\theta=133\degree"